wRITE THE EQUATION OF THE LINE L SATISFYING THE GIVEN GEOMETRIC CONDITIONS.
L has x-intercept (4,0) and Y intercept(0,2)

Answer by nerdybill(7384)

The "slope-intercept" form of a line is:
y = mx + b
where
m is the slope
b is the y-intercept at (0,b)
.
For any two points (on a straight line)
(x1,y1) and (x2,y2)
the slope is
m = (y2-y1)/(x2-x1)
.
In your problem, the two points were:
(4,0) and (0,2)
Substituting into:
m = (y2-y1)/(x2-x1)
m = (2-0)/(0-4)
m = (2)/(-4)
m = -1/2
.
The problem gave the y-intercept at (0,2)
Therefore, b = 2
.
Plugging it all back into:
y = mx + b
we have:
y = (-1/2)x + 2
or
y = -x/2 + 2